OpenStudy (anonymous):

4x-3y=3 find slope and explain

OpenStudy (anonymous):

4x-3y=3 or 3y =4x -3 or y=(4/3) x -1 (it is of the form y=mx+c) hence slope is 4/3

OpenStudy (anonymous):

Convert this into y = mx + b form first...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Come on go ahead..

OpenStudy (anonymous):

3y=-4x-3

OpenStudy (anonymous):

See : \(4x - 3y = 3\) Here : Add 3y both the sides first and tell me what did you get ??

OpenStudy (anonymous):

4x-6y=3

OpenStudy (anonymous):

I said to add 3y...

OpenStudy (anonymous):

oh

OpenStudy (anonymous):